# Inferring Outcome Means of Exponential Family Distributions Estimated by Deep Neural Networks

### v2

@@ 4 Numerical Experiments
$n=400,r=n^{0.8}$0.13 0.34(0.34) -0.33 0.45(0.52) 0.48 0.42 93.5% 1.82(0.75)
$r=n^{0.85}$0.07 0.32(0.33) -0.26 0.42(0.48) 0.46 0.44 95.8% 2.09(0.87)
$r=n^{0.9}$0.02 0.31(0.33) -0.20 0.39(0.44) 0.46 0.49 98.1% 2.50(1.08)
Boostrap with $f_{0}$0.20 0.76(0.80) 0.05 0.82(0.77) 1.10 - 95.1% 19.9(15.8)
$n=700,r=n^{0.8}$0.06 0.33(0.34) -0.29 0.43(0.49) 0.47 0.39 92.8% 1.83(0.78)
$r=n^{0.85}$0.01 0.31(0.34) -0.22 0.40(0.45) 0.46 0.41 95.3% 2.07(0.91)
$r=n^{0.9}$ 0.04 0.30(0.34) -0.16 0.37(0.41) 0.46 0.45 97.3% 2.44(1.13)
Boostrap with $f_{0}$0.24 0.70(0.70) -0.19 0.75(0.76) 0.99 - 94.7% 12.9(10.5)
Binomial Model
$n=400,r=n^{0.8}$ 0.29 0.63(0.59) 0.05 0.09(0.10) 0.86 0.60 86.5% 0.31(0.01)
$r=n^{0.85}$ 0.23 0.60(0.56) 0.05 0.08(0.09) 0.82 0.64 89.2% 0.33(0.03)
$r=n^{0.9}$ 0.19 0.57(0.52) 0.04 0.08(0.09) 0.77 0.71 93.0% 0.37(0.06)
Boostrap with $f_{0}$ 0.72 1.17(0.97) 0.06 0.13(0.12) 1.52 - 92.2% 0.55(0.28)
$n=700,r=n^{0.8}$ 0.07 0.49(0.45) 0.03 0.06(0.06) 0.67 0.55 92.5% 0.30(0.05)
$r=n^{0.85}$ 0.03 0.48(0.44) 0.02 0.06(0.06) 0.65 0.60 94.4% 0.33(0.06)
$r=n^{0.9}$0.01 0.47(0.44) 0.02 0.06(0.05) 0.65 0.68 96.2% 0.38(0.09)
Boostrap with $f_{0}$ 0.13 0.97(0.84) -0.01 0.11(0.11) 1.28 - 94.8% 0.61(0.27)
Aim (iv) investigates whether narrower yet deeper architectures can maintain robust performance under our Ensemble Subsampling Method (ESM). To explore this, we design a deeper network with $L=5$ and the layer structure $\mathbf{p}=(p,10,15,20,25,1)$. It is well known that training deep networks poses the risk of gradient explosion or vanishing during backpropagation (Glorot and Bengio, 2010). To improve training stability, we incorporate batch normalization layers after the second and second-to-last layers. These modifications aim to mitigate gradient instability while maintaining the network’s ability to learn complex representations under the ESM framework. Table 4 shows the performance of deeper, narrower neural networks. Despite the increased depth, prediction deviations remain small and comparable to those of shallow, wide networks. Key metrics, including empSD, SE and CP, align well with theoretical expectations, indicating robust inferential validity. However, AIL for deeper networks is somewhat wider than that in the baseline setting, as additional layers introduce non-linear interactions that may amplify variability in uncertainty estimates.
5 Analysis of eICU data
Predicting ICU readmission is vital for improving patient outcomes and reducing healthcare costs. Readmitted patients face higher mortality risks and increased expenses, while hospitals must optimize resource use. Identifying high-risk individuals allows for more personalized care, such as closer monitoring. The eICU Collaborative Research Database (eICU-CRD) (Pollard et al., 2018) is a large-scale, multi-center collection of anonymized ICU patient data, created through collaboration between the MIT Laboratory for Computational Physiology and healthcare organizations. It includes detailed information on patients’ physiological measurements, clinical data, lab results, medication usage, and outcomes during their ICU stays. Our analysis focuses on Hospital #188, a major hospital with 2,632 patients admitted to ICU. We focus on 20 laboratory results, including chloride, white blood cell (WBC) count, hemoglobin (Hgb), red cell distribution width (RDW), calcium, platelet count, mean corpuscular volume (MCV), and mean corpuscular hemoglobin concentration (MCHC), which have been identified in the medical literature as predictors of hospitalization. Combined with other patient demographic information, this results in a total of 55 covariates for analysis. Regarding outcomes, among the 2,632 patients admitted to the ICU, 1,476 were not readmitted, 1,046 were readmitted once, 77 were readmitted twice, 26 were readmitted three times, and 7 were readmitted five or more times. We apply nonparametric logistic and Poisson models to, respectively, predict ICU readmission and the total number of ICU readmissions using our ESM. To evaluate the effectiveness of our methods, we perform 5-fold cross-validation. Specifically, we divide the data into five equal-sized subsamples, train the model on four subsamples, and use the ESM method to estimate the mean and corresponding inference interval for the held-out fold. This process is repeated across all five folds, allowing us to obtain point estimates and intervals for all sample points.
Figure 3: Evaluation in the logistic model. Figure 3(a) presents the ROC curve for the logistic model, illustrating the model’s ability to estimate prediction mean, with an AUC of 0.86. Figure 3(b) displays individual estimates along with their corresponding confidence intervals, highlighting the heteroskedasticity observed across individuals.
Figure 4: Histogram under distinct probability intervals in logistic models. Figures 4(a) and 4(b) present hemoglobin (Hbg) and red cell distribution width (RDW) under different probability intervals. The blue histograms represent patients for whom the whole probability interval is above 0.7, whereas the orange histograms represent patients for whom the whole probability interval is below 0.15.
For predicting ICU readmission, we define the binary outcome as: $y=0$ for readmission times of zero, and $y=1$ for readmission times greater than or equal to one. A high probability of $y=1$ predicted based on covariates suggests a high likelihood of ICU readmission for a given patient profile, indicating the need for more personalized care for these patients. To implement our method, we set the subsampling size to $r=n^{0.9}=765$ and the number of subsampling iterations to $B=3000$. The neural network used has the structure $(55,128,64,1)$ with the ReLU activation function. During training, we fine-tune the learning rate to $\eta=0.2$, use 500 training epochs, and set the weight decay parameter to 0.03.
As reported in Figure 3(a), the Area under the Receiver Operating Characteristic Curve (AUC-ROC) is around $0.86$, showing a satisfactory performance of using DNN-estimated $\mathbb{E}(y|\mathbf{x})$ to distinguish patients. Further, Figure 3(b) shows the estimated probabilities and their corresponding confidence intervals across individuals. The $x$-axis represents individual cases (sorted by estimated probability for clarity), and the $y$-axis shows the estimated probability. The gray intervals indicate uncertainty, constructed using a significance level of $\alpha=0.05$. Notably, both the estimated probabilities and interval lengths vary across individuals, reflecting heteroskedasticity in the model’s predictions. Even among individuals with similar predicted probabilities, uncertainty can differ, suggesting differences in data quality or individual characteristics that make some patients inherently more or less predictable.
Moreover, some interesting observations are available. Figure 4(a) shows patients with probability interval above 0.7 of ICU readmission exhibit significantly lower hemoglobin levels compared to those with probability interval below 0.3. Since hemoglobin is critical for oxygen transport, reduced levels may reflect compromised tissue oxygenation, potentially exacerbating organ dysfunction and increasing the likelihood of clinical deterioration that necessitates readmission. Additionally, Figure 4(b) shows that higher RDW (red cell distribution width) is also observed among patients with ICU readmission. Higher RDW is often associated with systemic inflammation, nutritional deficiencies or impaired red blood cell production, all of which could signal unresolved physiologic stress or comorbidities that delay recovery and predispose patients to repeated critical care interventions, which may be factors that could contribute to the increased risk of ICU readmission. These findings align with previous studies (Vincent et al., 2002; Lippi and Plebani, 2014), further supporting the role of hematologic markers in predicting ICU readmission.
We also fit a nonparametric Poisson regression model by using DNN to estimate $\mathbb{E}(y|\mathbf{x}_{*})$ for a new data point $\mathbf{x}_{*}$. The results may provide further insights into patient well-being with more detailed count data. Based on our numerical experience, we find that the same hyperparameters used in the nonparametric logistic regression model perform well in this case as well.
[…]

### v3

@@ A.4 Additional Experiments on Dimensionality and Nonlinearity
$n=400,r=n^{0.8}$ -0.42(0.50) 0.50(0.42) -0.23(0.34) 0.29(0.29) 0.42 0.52 0.44 85.6% 0.97(0.66)
$r=n^{0.85}$ -0.29(0.48) 0.42(0.36) -0.16(0.33) 0.26(0.26) 0.41 0.52 0.43 90.3% 1.09(0.76)
$r=n^{0.9}$ -0.18(0.45) 0.37(0.32) -0.10(0.35) 0.25(0.27) 0.40 0.51 0.42 92.7% 1.22(0.93)
$n=700,r=n^{0.8}$ -0.21(0.37) 0.32(0.28) -0.14(0.27) 0.21(0.22) 0.30 0.41 0.31 88.0% 0.78(0.52)
$r=n^{0.85}$ -0.12(0.35) 0.28(0.24) -0.08(0.28) 0.20(0.21) 0.29 0.40 0.30 90.5% 0.83(0.57)
$r=n^{0.9}$ -0.05(0.34) 0.26(0.22) -0.05(0.27) 0.19(0.20) 0.27 0.39 0.28 91.2% 0.84(0.58)
As shown in Table S.7, increasing the dimension to $p=20$ degrades estimation accuracy, with $\text{MAE}_{f}$ rising from 0.28 to 0.55 and empirical standard deviation from 0.34 to 0.62 under the logistic regression model with $n=700$. This reflects increased bias and variability, leading to a drop in coverage from 93.7% to 92.1%. Similar effects of high-dimensional noise on random forests were observed in Wager & Athey (2018). Increasing the sample size to $n=2000$ under $p=20$ improves estimation, restoring coverage to nearly 95%. Conversely, reducing the dimension to $p=7$ improves neural network performance (Table S.8), with $\text{MAE}_{f}$ decreasing to 0.23 and empirical standard deviation to 0.27. The reduced bias yields shorter intervals and coverage close to 95%.
We also consider another nonlinear function
$g(\mathbf{x})=2\tanh\bigg(\frac{1.5x_{1}+0.6(x_{2}^{2}-1)+0.4x_{3}\cdot\tanh(x_{4})+0.15\cdot\sin(x_{5})}{2.5}\bigg),$
where $\tanh(x)=(e^{x}-e^{-x})/(e^{x}+e^{-x})$. This function introduces quadratic, $\tanh$, and even interaction components, representing a moderately nonlinear structure that remains learnable for neural networks. All other simulation settings are kept the same as the basic settings as those in Section 4.
Table S.9 shows the results with a different function $g(\mathbf{x})$. We find that the corrected standard error estimator remains accurate, but the coverage probability shows a mild drop compared to the basic setting. This phenomenon shows that not only random forests but also neural networks may exhibit bias in their output logits, particularly under complex nonlinear structures. Neural networks, especially in nonlinear and high-dimensional regimes, can incur non-negligible approximation bias due to optimization challenges or limitations in representing the true underlying function $g(\mathbf{x})$. Such biases are difficult to eliminate in practice and can result in reduced coverage probability, even when variance estimation is accurate. How to effectively reduce or correct for such bias remains an open problem, and addressing it is key to improving the reliability of inference.
A.5 Additional Experiments with Optimization
We examine how optimization hyperparameters, i.e., weight decay and training epochs, affect results and find our inference stable across reasonable settings. In particular, we vary training epochs (300, 700) and weight decay ($2\times 10^{-3}$, $2\times 10^{-4}$).
Table S.10: Simulation results with different epochs.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
Logistic with Epochs =300=300
$n=400,r=n^{0.8}$ 0.08(0.50) 0.40(0.31) 0.01(0.10) 0.08(0.06) 0.46 0.56 0.47 93.4% 0.36(0.09)
$r=n^{0.85}$ 0.07(0.44) 0.35(0.28) 0.02(0.09) 0.07(0.06) 0.41 0.50 0.42 93.9% 0.34(0.08)
$r=n^{0.9}$ 0.07(0.40) 0.32(0.26) 0.01(0.09) 0.07(0.06) 0.38 0.46 0.39 93.9% 0.32(0.07)
$n=700,r=n^{0.8}$ 0.07(0.33) 0.26(0.21) 0.01(0.07) 0.06(0.05) 0.29 0.39 0.30 92.5% 0.25(0.05)
$r=n^{0.85}$ 0.07(0.31) 0.25(0.19) 0.02(0.07) 0.05(0.04) 0.27 0.36 0.27 91.6% 0.23(0.05)
$r=n^{0.9}$ 0.06(0.30) 0.24(0.19) 0.01(0.07) 0.05(0.04) 0.25 0.35 0.26 90.0% 0.22(0.05)
Logistic with Epochs =700=700
$n=400,r=n^{0.8}$ 0.13(0.86) 0.69(0.53) 0.02(0.15) 0.13(0.09) 0.76 0.93 0.78 90.5% 0.51(0.15)
$r=n^{0.85}$ 0.08(0.72) 0.57(0.45) 0.01(0.14) 0.11(0.08) 0.67 0.83 0.70 92.8% 0.49(0.14)
$r=n^{0.9}$ 0.06(0.61) 0.48(0.38) 0.01(0.12) 0.10(0.07) 0.58 0.72 0.60 94.1% 0.45(0.12)
$n=700,r=n^{0.8}$ 0.07(0.50) 0.40(0.31) 0.01(0.10) 0.08(0.06) 0.46 0.62 0.47 92.9% 0.36(0.09)
$r=n^{0.85}$ 0.05(0.41) 0.32(0.25) 0.01(0.09) 0.07(0.05) 0.39 0.52 0.40 94.5% 0.32(0.08)
$r=n^{0.9}$ 0.05(0.36) 0.28(0.22) 0.01(0.08) 0.06(0.05) 0.34 0.46 0.35 94.5% 0.29(0.07)
Poisson with Epochs =300=300
$n=400,r=n^{0.8}$ -0.23(0.32) 0.30(0.25) -0.13(0.22) 0.19(0.16) 0.30 0.37 0.31 90.2% 0.78(0.40)
$r=n^{0.85}$ -0.15(0.31) 0.27(0.22) -0.08(0.23) 0.18(0.16) 0.30 0.37 0.30 92.7% 0.84(0.45)
$r=n^{0.9}$ -0.08(0.30) 0.24(0.20) -0.04(0.23) 0.17(0.16) 0.29 0.36 0.30 94.2% 0.88(0.48)
$n=700,r=n^{0.8}$ -0.12(0.23) 0.20(0.16) -0.07(0.17) 0.14(0.13) 0.21 0.29 0.22 92.2% 0.60(0.30)
$r=n^{0.85}$ -0.06(0.22) 0.18(0.14) -0.04(0.17) 0.13(0.12) 0.21 0.29 0.21 93.8% 0.62(0.31)
$r=n^{0.9}$ -0.01(0.22) 0.17(0.14) -0.01(0.17) 0.13(0.12) 0.20 0.28 0.20 93.7% 0.62(0.31)
Poisson with Epochs =700=700
$n=400,r=n^{0.8}$ -0.37(0.41) 0.43(0.34) -0.19(0.24) 0.24(0.19) 0.37 0.47 0.39 87.2% 0.91(0.49)
$r=n^{0.85}$ -0.26(0.39) 0.36(0.30) -0.13(0.26) 0.22(0.19) 0.37 0.47 0.39 92.0% 1.02(0.57)
$r=n^{0.9}$ -0.18(0.37) 0.32(0.26) -0.08(0.26) 0.21(0.18) 0.35 0.46 0.38 93.9% 1.10(0.65)
$n=700,r=n^{0.8}$ -0.20(0.29) 0.27(0.22) -0.11(0.20) 0.17(0.15) 0.27 0.37 0.28 90.3% 0.72(0.37)
$r=n^{0.85}$ -0.12(0.27) 0.23(0.19) -0.07(0.20) 0.16(0.14) 0.26 0.36 0.27 93.1% 0.76(0.40)
$r=n^{0.9}$ -0.06(0.26) 0.21(0.17) -0.03(0.21) 0.15(0.14) 0.25 0.35 0.26 94.3% 0.77(0.42)
Table S.11: Simulation results with different weight decay.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
Logistic with WeightDecay $2\times 10^{-3}$
$n=400,r=n^{0.8}$ 0.18(1.61) 1.28(0.98) 0.02(0.22) 0.19(0.11) 1.25 1.54 1.27 87.5% 0.62(0.22)
$r=n^{0.85}$ 0.13(1.30) 1.02(0.81) 0.02(0.20) 0.16(0.10) 1.08 1.36 1.13 90.9% 0.61(0.19)
$r=n^{0.9}$ 0.08(1.06) 0.83(0.66) 0.01(0.18) 0.15(0.10) 0.94 1.19 0.97 92.9% 0.59(0.17)
$n=700,r=n^{0.8}$ 0.10(0.93) 0.73(0.58) 0.01(0.15) 0.13(0.08) 0.75 1.02 0.76 89.2% 0.48(0.15)
$r=n^{0.85}$ 0.07(0.72) 0.57(0.45) 0.01(0.13) 0.11(0.08) 0.63 0.87 0.64 92.0% 0.45(0.12)
$r=n^{0.9}$ 0.03(0.59) 0.46(0.37) 0.00(0.11) 0.09(0.07) 0.54 0.76 0.55 93.8% 0.41(0.10)
Logistic with WeightDecay $2\times 10^{-4}$
$n=400,r=n^{0.8}$ 0.22(1.72) 1.36(1.07) 0.02(0.22) 0.19(0.11) 1.32 1.66 1.37 88.1% 0.64(0.23)
$r=n^{0.85}$ 0.15(1.41) 1.11(0.88) 0.02(0.20) 0.17(0.11) 1.17 1.48 1.22 91.1% 0.64(0.20)
$r=n^{0.9}$ 0.09(1.13) 0.88(0.71) 0.01(0.18) 0.15(0.10) 1.00 1.28 1.04 93.3% 0.61(0.17)
$n=700,r=n^{0.8}$ 0.12(1.00) 0.80(0.63) 0.02(0.16) 0.14(0.09) 0.80 1.10 0.82 89.1% 0.50(0.16)
$r=n^{0.85}$ 0.07(0.78) 0.61(0.49) 0.01(0.14) 0.11(0.08) 0.67 0.93 0.69 92.0% 0.47(0.13)
$r=n^{0.9}$ 0.04(0.64) 0.50(0.40) 0.01(0.12) 0.10(0.07) 0.59 0.83 0.60 93.6% 0.43(0.11)
Poisson with WeightDecay $2\times 10^{-3}$
$n=400,r=n^{0.8}$ -0.67(0.60) 0.70(0.56) -0.28(0.25) 0.31(0.21) 0.49 0.63 0.51 83.0% 0.95(0.51)
$r=n^{0.85}$ -0.49(0.56) 0.56(0.49) -0.21(0.27) 0.27(0.20) 0.48 0.62 0.50 88.8% 1.10(0.60)
$r=n^{0.9}$ -0.35(0.50) 0.46(0.41) -0.14(0.29) 0.25(0.20) 0.46 0.61 0.49 92.9% 1.24(0.72)
$n=700,r=n^{0.8}$ -0.41(0.41) 0.44(0.37) -0.20(0.21) 0.23(0.17) 0.35 0.49 0.36 84.3% 0.77(0.39)
$r=n^{0.85}$ -0.27(0.37) 0.35(0.30) -0.13(0.22) 0.20(0.16) 0.33 0.48 0.34 90.5% 0.84(0.43)
$r=n^{0.9}$ -0.18(0.35) 0.30(0.26) -0.08(0.24) 0.19(0.16) 0.33 0.48 0.33 92.8% 0.90(0.49)
Poisson with WeightDecay $2\times 10^{-4}$
$n=400,r=n^{0.8}$ -0.72(0.64) 0.75(0.61) -0.29(0.25) 0.32(0.22) 0.52 0.65 0.53 81.6% 0.96(0.51)
$r=n^{0.85}$ -0.54(0.59) 0.60(0.53) -0.22(0.27) 0.29(0.21) 0.51 0.65 0.52 88.4% 1.11(0.61)
$r=n^{0.9}$ -0.38(0.52) 0.48(0.42) -0.16(0.29) 0.26(0.21) 0.47 0.64 0.51 92.7% 1.26(0.73)
$n=700,r=n^{0.8}$ -0.44(0.43) 0.48(0.39) -0.21(0.21) 0.24(0.17) 0.36 0.51 0.37 83.1% 0.77(0.39)
$r=n^{0.85}$ -0.30(0.39) 0.38(0.33) -0.14(0.23) 0.21(0.17) 0.35 0.50 0.36 89.6% 0.85(0.44)
$r=n^{0.9}$ -0.20(0.36) 0.31(0.27) -0.09(0.24) 0.19(0.16) 0.33 0.50 0.34 92.9% 0.92(0.50)
As shown in Tables S.10 and S.11, our variance estimator remains accurate across different training epochs and weight decay values, closely matching the empirical standard deviation. This confirms that optimization error is negligible under reasonable hyperparameter settings and supports viewing Assumption 3.3 as a technical device separating optimization from inference rather than a practical constraint on SGD convergence. Larger subsample sizes $r$ reduce bias in MAE, consistent with modern generalization theory. Overall, the results validate the robustness of our inference guarantees, while highlighting that extreme hyperparameters (e.g., excessive weight decay or too few epochs) can impair learning and should be avoided.
A.6 Additional Experiments on $r$
Recall that $r=n^{\gamma}$. We vary $\gamma$ from 0.65 to 0.95 (step 0.05) while keeping other settings as in Section 4, and plot the standard deviation, MAE of $\widehat{f}^{B}-f_{0}$, and coverage probability versus $\gamma$ for $n=400$ and $n=700$ under two models.
Figure S.2: Simulation results with different values of $\gamma$. Blue and red lines/points correspond to $n=400$ and $n=700$ respectively. Figures 2(a)2(c) present the results under the logistic model, and Figures 2(d)2(f) show results under the Poisson model.
Figures 2(a)2(d) show that as $\gamma$ increases, our IJ variance estimator closely matches the empirical standard deviation in both logistic and Poisson models, confirming its robustness and validating the finite-sample correction factor $n(n-1)/(n-r)^{2}$. Figures 2(b)2(e) indicate that MAE decreases with larger $r$, supporting Theorem 3.5 and showing bias reduction with larger subsamples. Finally, Figures 2(c)2(f) demonstrate that as $r$ increases, coverage improves toward the nominal 95% level, consistent with our assumption on $\gamma$ ensuring bias decay and valid inference.
We also evaluate computational cost across different $r$ values by recording the training time of a single neural network. All the simulations are conducted in the institute high performance computing platform. The simulation is repeated 300 times, with $B$ networks trained per repetition. As shown in Table S.12, training one network takes only a few seconds, while the main cost arises from repeatedly training the $B$ networks. Following Wager & Athey (2018) and Fei & Li (2024), we let $B$ grow proportionally with $n$ and apply bias correction to maintain variance estimation accuracy, largely reducing computation. Although $300\times B$ trainings would otherwise be intensive, high-performance computing with parallel nodes and multi-CPU systems makes the experiments feasible.
Table S.12: Computation time in seconds.
$\gamma=$ $0.65$ $0.7$ $0.75$ $0.8$ $0.85$ $0.9$ $0.95$
Logistic 0.74 0.82 1.00 1.38 1.41 1.90 2.30
Poisson 0.75 0.90 0.96 1.32 1.41 1.78 2.24
A.7 Additional Experiments on $B$
We decrease $B$ from 1400 to 400 in increments of 200 to assess the consistency of variance estimation, while fixing $\gamma=0.9$.
Figure S.3: Simulation results with different values of $B$. Blue and red lines/points correspond to $n=400$ and $n=700$ respectively. Figures 3(a)3(c) present the results under the logistic model, and Figures 3(d)3(f) show results under the Poisson model.
Figure S.3 reports the standard deviations, $\text{MAE}_{f}$, and coverage probabilities as $B$ varies, analogous to Section A.6. As $B$ increases, the gap between empSD and the uncorrected SE narrows, while our bias-corrected estimator $\text{SE}_{c}$ closely matches empSD even when $B=400$ and $n=700$ ($B<n$). The $\text{MAE}_{f}$ is smaller for $n=700$ than for $n=400$, and all coverage probabilities remain near the nominal 95% level. These results show that the bias correction is effective even with smaller $B$, allowing valid inference and stable performance while reducing computational cost.
Table S.13: Simulation results for different models under varying sample sizes $n$ and subsampling ratios $r$ with larger $B=3000$.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
Logistic Model
$n=400,r=n^{0.8}$ 0.10(0.71) 0.56(0.44) 0.02(0.13) 0.11(0.08) 0.64 0.72 0.66 92.0% 0.46(0.13)
$r=n^{0.85}$ 0.07(0.60) 0.47(0.37) 0.01(0.12) 0.10(0.07) 0.56 0.63 0.57 93.4% 0.43(0.11)
$r=n^{0.9}$ 0.05(0.49) 0.39(0.31) 0.01(0.10) 0.08(0.06) 0.47 0.55 0.50 95.3% 0.39(0.09)
$n=700,r=n^{0.8}$ 0.06(0.41) 0.33(0.26) 0.01(0.09) 0.07(0.05) 0.38 0.46 0.40 93.7% 0.31(0.08)
$r=n^{0.85}$ 0.05(0.35) 0.28(0.22) 0.01(0.08) 0.06(0.05) 0.33 0.39 0.34 94.2% 0.28(0.06)
$r=n^{0.9}$ 0.05(0.33) 0.26(0.20) 0.01(0.07) 0.06(0.04) 0.30 0.36 0.31 93.8% 0.26(0.06)
Poisson Model
$n=400,r=n^{0.8}$ -0.32(0.38) 0.39(0.32) -0.17(0.24) 0.22(0.18) 0.36 0.39 0.36 88.1% 0.87(0.46)
$r=n^{0.85}$ -0.21(0.37) 0.32(0.27) -0.10(0.25) 0.20(0.18) 0.34 0.39 0.36 92.2% 0.96(0.53)
$r=n^{0.9}$ -0.13(0.35) 0.28(0.24) -0.06(0.26) 0.20(0.18) 0.33 0.38 0.35 93.5% 1.02(0.58)
$n=700,r=n^{0.8}$ -0.16(0.27) 0.24(0.20) -0.09(0.19) 0.16(0.14) 0.25 0.30 0.25 90.4% 0.68(0.34)
$r=n^{0.85}$ -0.10(0.26) 0.21(0.17) -0.06(0.19) 0.15(0.14) 0.24 0.30 0.25 93.1% 0.71(0.37)
$r=n^{0.9}$ -0.04(0.24) 0.19(0.15) -0.02(0.19) 0.14(0.13) 0.23 0.28 0.24 95.2% 0.72(0.38)
Figure S.4: Estimation and inference in simulation samples: Logistic Model with $n=700$, $r=n^{0.9}$ and $B=3000$. Figure 4(a) shows the average estimated $\mathbb{E}(y|\mathbf{x})$ with variability across 300 repetitions (gray band) over test points. Figure 4(b) compares corrected and uncorrected standard errors and their variability (gray and blue bands, respectively) to the empirical standard deviations of estimates across all test samples.
We further assess the impact of a larger subsample size ($B=3000$) on variance estimation and the accuracy of the correction term. Table S.13 shows that the results remain consistent: increasing $B$ reduces MAE for both $\widehat{f}^{B}$ and $\psi^{\prime}(\widehat{f}^{B})$, and the empirical standard deviations (empSD) align closely with the estimated standard errors (SE), maintaining nominal 95% coverage. Figures 4(a) and 4(b) illustrate that our estimator accurately recovers the conditional mean and that the corrected variance closely follows empSD, whereas the uncorrected version slightly overestimates. As $B$ increases from 1400 to 3000, this bias diminishes, consistent with theory showing that the correction term vanishes asymptotically. These findings confirm that the correction improves finite-sample accuracy while preserving asymptotic validity.
A.8 Additional Experiments on $L$
We now examine how network depth affects estimation bias and variance using $n=700$, varying both $L$ and $r$. For each $L$, we set $\mathbf{p}=(p,\underbrace{30,\ldots,30}_{L-1},1)$ and apply batch normalization after the 2nd, 4th, and 6th layers to ensure gradient stability. We consider $L=3,4,5,6$ and $r=n^{0.8},n^{0.85},n^{0.9},n^{0.95}$.
Tables S.14S.15 show that deeper networks tend to introduce larger biases, consistent with (C.2), which predicts larger approximation error and lower coverage for deeper architectures. Hence, in our settings, it appears that relatively shallow networks are preferable for both stability and reduced bias (Goodfellow et al. 2016).
Consistent with the sensitivity analysis in Section A.6, increasing $r$ reduces $\text{MAE}_{f}$, and all bias-corrected variance estimators closely match the empirical results, confirming the robustness of our inference procedure.
Table S.14: Simulation results with Logistic under varying $L$ and $r$.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
Logistic
$r=n^{0.8},L=3$ 0.19(0.66) 0.57(0.39) 0.04(0.13) 0.11(0.08) 0.59 0.92 0.60 84.5% 0.43(0.15)
$L=4$ 0.21(0.71) 0.61(0.41) 0.04(0.14) 0.12(0.08) 0.61 0.98 0.61 82.7% 0.44(0.16)
$L=5$ 0.21(0.73) 0.63(0.43) 0.04(0.14) 0.12(0.08) 0.63 1.11 0.63 82.0% 0.44(0.17)
$L=6$ 0.21(0.70) 0.61(0.41) 0.04(0.14) 0.12(0.08) 0.61 1.10 0.63 83.4% 0.44(0.16)
$r=n^{0.85},L=3$ 0.19(0.69) 0.59(0.41) 0.04(0.14) 0.12(0.08) 0.63 1.03 0.64 85.6% 0.46(0.16)
$L=4$ 0.21(0.74) 0.64(0.43) 0.04(0.14) 0.13(0.08) 0.66 1.12 0.67 83.2% 0.46(0.17)
$L=5$ 0.20(0.80) 0.69(0.46) 0.04(0.15) 0.13(0.08) 0.71 1.32 0.71 81.8% 0.48(0.19)
$L=6$ 0.21(0.77) 0.66(0.45) 0.04(0.15) 0.13(0.09) 0.69 1.31 0.70 83.0% 0.48(0.18)
$r=n^{0.9},L=3$ 0.19(0.71) 0.60(0.42) 0.04(0.14) 0.12(0.08) 0.65 1.16 0.68 86.8% 0.48(0.16)
$L=4$ 0.21(0.75) 0.64(0.44) 0.04(0.15) 0.13(0.08) 0.69 1.27 0.71 84.9% 0.49(0.18)
$L=5$ 0.21(0.85) 0.73(0.48) 0.04(0.16) 0.14(0.09) 0.76 1.60 0.78 81.9% 0.51(0.21)
$L=6$ 0.20(0.82) 0.72(0.47) 0.04(0.16) 0.14(0.09) 0.75 1.57 0.76 82.7% 0.51(0.20)
$r=n^{0.95},L=3$ 0.17(0.71) 0.59(0.42) 0.03(0.14) 0.12(0.08) 0.66 1.37 0.69 88.9% 0.49(0.16)
$L=4$ 0.20(0.74) 0.63(0.44) 0.04(0.15) 0.12(0.09) 0.69 1.48 0.72 86.5% 0.49(0.18)
$L=5$ 0.19(0.90) 0.77(0.50) 0.03(0.17) 0.15(0.09) 0.82 2.14 0.84 81.7% 0.54(0.23)
$L=6$ 0.21(0.86) 0.73(0.49) 0.04(0.17) 0.14(0.09) 0.79 2.07 0.82 81.9% 0.53(0.23)
Table S.15: Simulation results with Poisson under varying $L$ and $r$.
$\text{Bias}_{f}$ $\text{MAE}_{f}$ $\text{Bias}_{\psi^{\prime}}$ $\text{MAE}_{\psi^{\prime}}$ EmpSD SE $\text{SE}_{c}$ CP AIL
Poisson
$r=n^{0.8},L=3$ -0.38(0.33) 0.41(0.29) -0.20(0.21) 0.23(0.17) 0.31 0.49 0.31 79.8% 0.71(0.37)
$L=4$ -0.42(0.37) 0.45(0.32) -0.20(0.21) 0.24(0.17) 0.33 0.53 0.33 76.9% 0.74(0.40)
$L=5$ -0.47(0.38) 0.50(0.34) -0.23(0.22) 0.26(0.17) 0.34 0.59 0.34 72.9% 0.75(0.44)
$L=6$ -0.48(0.39) 0.51(0.35) -0.22(0.22) 0.26(0.17) 0.34 0.62 0.35 73.7% 0.78(0.46)
$r=n^{0.85},L=3$ -0.30(0.35) 0.36(0.28) -0.15(0.22) 0.21(0.17) 0.33 0.55 0.33 87.0% 0.82(0.43)
$L=4$ -0.35(0.38) 0.41(0.31) -0.16(0.23) 0.22(0.17) 0.35 0.61 0.36 84.6% 0.86(0.46)
$L=5$ -0.42(0.41) 0.48(0.34) -0.19(0.24) 0.25(0.18) 0.37 0.69 0.37 79.6% 0.88(0.52)
$L=6$ -0.43(0.42) 0.48(0.35) -0.19(0.24) 0.25(0.18) 0.38 0.72 0.38 80.0% 0.90(0.53)
$r=n^{0.9},L=3$ -0.22(0.36) 0.33(0.27) -0.10(0.24) 0.20(0.17) 0.34 0.63 0.35 91.2% 0.94(0.50)
$L=4$ -0.26(0.39) 0.37(0.30) -0.11(0.25) 0.21(0.17) 0.37 0.69 0.38 89.1% 0.98(0.52)
$L=5$ -0.35(0.43) 0.44(0.34) -0.15(0.26) 0.24(0.18) 0.40 0.82 0.40 84.8% 1.02(0.60)
$L=6$ -0.36(0.45) 0.45(0.35) -0.14(0.27) 0.25(0.18) 0.41 0.86 0.41 84.5% 1.05(0.62)
$r=n^{0.95},L=3$ -0.14(0.37) 0.30(0.25) -0.04(0.27) 0.20(0.18) 0.35 0.77 0.37 93.1% 1.06(0.62)
$L=4$ -0.17(0.40) 0.33(0.28) -0.05(0.27) 0.21(0.18) 0.37 0.84 0.39 91.6% 1.11(0.62)
$L=5$ -0.26(0.45) 0.41(0.32) -0.09(0.29) 0.24(0.19) 0.42 1.06 0.43 87.8% 1.20(0.74)
$L=6$ -0.27(0.46) 0.42(0.33) -0.09(0.29) 0.24(0.19) 0.42 1.11 0.44 87.2% 1.22(0.76)
A.9 Additional Experiments on Dropout Rate
We vary the dropout rates, which are set to 0.3 and 0.5 in the two models, while retaining all other settings as in Section 4.
Table S.16: Simulation results with different dropout rate.
[…]